Potassium nitrate has a lattice energy of -163.8 kcal/mol and a heat of hydration of -155.5 kcal/mol.

How much potassium nitrate has to dissolve in water to absorb 115kJ of heat?

Express your answer using two significant figures. (in kg)
Answer #1

enthalpy of solution = lattice energy + hydration energy

let x mole of KNO3 be dissolved

so

115 KJ = 27.486 Kcal

27.486 = x(163.8 - 155.5)

x = 3.312 moles

mass = 3.312*166 = 549.72 gm = 0.549 Kg
Answer #2

lattice energy + hydration energy + enthalpy of solution = 0

let x mole of KNO3 be dissolved

1 kcal = 4.186 kJ

115 KJ = 27.486 Kcal

x(-163.8 - 155.5) + 27.486 = 0

x = 27.486 / ( 163.8 + 155.5 ) = 0.086 moles

1 mole = 166 kg

mass = 0.086*166 = 14.276 gm = 0.014276 kg = 0.014 kg
Answer #3

?H(soln) = -U(KNO3) + ?Hhyd(KNO3) U is the lattice energy energy added to break up lattice so +ve
Convert to kJ: U = -163.8
Answer #4

_____Summing dHs

____KNO3 (s) ----___--> KNO3 (bits) -___----> solvated KNO3 (bits)
____________-163.8kcal/mole_________dH